OpenStudy (anonymous):

Which ordered pair is NOT a solution of the inequality, 3x - 2y 12? Answer (-3, -10.5) (-1, -7) (0, -6) (2, -3)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

need to know how to do it?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

you forgot > btw

OpenStudy (anonymous):

lol right, how did you know?

OpenStudy (anonymous):

seen this problem before

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

same exactly

OpenStudy (anonymous):

keystone :)

OpenStudy (anonymous):

to solve this you must plug in each of the ordered pairs into the inequality, 3x - 2y >12 A. 3(-3) - 2(-10.5) >12 = 12 B. 3(-1) - 2(-7) >12 = 11 C. 3(0) - 2(-6) >12 = 12 D.3(2) - 2(-3) >12 = 12

OpenStudy (anonymous):

uhu

OpenStudy (anonymous):

the answer is B since 3(-1) - 2(-7) >12 = 11 and 11 is not > 12

OpenStudy (anonymous):

ohh i see now

OpenStudy (anonymous):

thx